# A Population Analysis of 20 Exoplanets Observed from the Optical to the Near-infrared Wavelengths with HST: Evidence for Widespread Stellar Contamination

### v2

@@ F.1 GJ 436 b
Stellar Activity: Reported S-indexes () and values: a value of = 0.620 and a corresponding = -5.298 were reported in Knutson et al. (2010), Sreejith et al. (2020) report a = -5.090.001 and most recently, Claudi et al. (2024) report an S-index of = 0.7350.321 and a corresponding = -5.32630.0347. Considering the Ca H&K lines in isolation, all of the reported values in the literature are consistent and indicative of a lower level of stellar activity for GJ 436. Multiple observations spanning almost a full proposed stellar activity cycle (7.750.10 yrs), revealed large variations 38%3% in the summed flux of the major FUV emission lines, and a multitude of brief flares (Loyd et al., 2023). Spitzer Space Telescope showed variations in transit depth in the same bandpass across different visits, an indication for occultation of stellar active regions (Knutson et al., 2011). Although our Bayes factor results suggest a low preference for an active star model, the SAD favours a substantial degree of stellar contamination. However, caution in the interpretation of this metric is advised due to the considerable uncertainties associated with the optical data.
@@ F.2 GJ 3470 b
Stellar Activity: Reported S-indexes () and values: Sreejith et al. (2020) report a = -4.86. Claudi et al. (2024) report an S-index of = 1.7550.321 and a corresponding = -5.06250.0843. The inconsistency between these reported measurements could be due to differing activity levels at each epoch, although broadly speaking both of the above literature values are indicative of a low-to-moderate activity regime for GJ 3470. STIS observations as analysed by Benneke et al. (2019) do not display any signs of stellar activity, while our interpretation indicate a low level of activity across the stellar metrics we defined, with substantial spectral temporal variations. The monitoring of flaring activity with the Cosmic Origins Spectrograph (COS) provided the detection of three flares (Bourrier et al., 2021), on top of confirming the absorption signature originating from GJ 3470 b’s exosphere in the stellar Lyman- line.
@@ F.3 HAT-P-1 b
Stellar Activity: Reported S-indexes () and values: In the discovery paper, Bakos et al. (2007a) report a value of = -4.720.05. Knutson et al. (2010) calculate an S-index of = 0.158 and a corresponding = -4.984. Danielski et al. (2022) find an S-index of = 0.15 0.00 and a corresponding = -5.110.07, and Claudi et al. (2024) report an S-index of = 0.1510.014 and a corresponding = -5.04080.1151, or = -4.89640.1404 when the correction detailed in Mittag et al. (2013) is applied. The literature values are consistent and all imply a low-to-moderate level of stellar activity for HAT-P-1. Our Bayes factor results return two of the highest values across the whole sample, indicating a high level of stellar activity. However, the SAD and SAT indicate a low preference for the stellar contamination model and show only minor temporal deviations among the STIS G430L observations which were taken four months apart.
@@ F.4 HAT-P-12 b
Stellar Activity: Reported S-indexes () and values: Knutson et al. (2010) report an S-index of = 0.253 and a corresponding = -5.104. Observations from two different epochs were analysed by Mancini et al. (2018) yielding values of = 0.3640.054 and corresponding = -4.880.07 (14.03.2015) and = 0.3750.075 and corresponding = -4.870.09 (24.04.2015) respectively. Claudi et al. (2024) find an S-index of = 0.3620.063 and a corresponding = -4.91480.0789, or = -4.53830.0946 when the correction detailed in Mittag et al. (2013) is applied. Some variation is observed in the literature values but broadly they are suggestive of low-to-moderate activity for HAT-P-12. Jiang et al. (2021) report that the discrepant chemical results among different transmission studies can be attributed to varying levels of unocculted stellar spots and faculae. Using the GTC OSIRIS observatory centred in the optical (500-900 nm), they infer a photosphere covered by 33% of spots and 9% of faculae. We find an overall quiet star across all our metrics, even though the SAT reported is moderately high within our the sample at 6.407% +- 1.883%, indicating a relatively large temporal variability across 20 days during which the system was observed with STIS G430L.
@@ F.5 HAT-P-18 b
Stellar Activity: Reported S-indexes () and values: Claudi et al. (2024) report an S-index of = 0.3340.079 and corresponding = -4.85530.1295, or = -4.52820.164 when the correction detailed in Mittag et al. (2013) is applied. These values are consistent with a moderate-to-high level of activity for HAT-P-12. Recent JWST observations with NIRISS are also suggestive of the presence of stellar contamination (Fournier-Tondreau et al., 2024). Both the Bayes Factor and the SAD metric agree to a moderate-to-high level of stellar contamination.
@@ F.6 HAT-P-26 b
Stellar Activity: Reported S-indexes () and values: In the discovery paper, Hartman et al. (2011a) report an S-index of = 0.182 and a corresponding = -4.992. Claudi et al. (2024) determine an S-index of = 0.1720.079 and a corresponding = -5.03720.2043, or = -4.85280.2747 when the correction detailed in Mittag et al. (2013) is applied. Literature values are consistent with a lower level of stellar activity for HAT-P-26. The photometric follow-up of the host star did not reveal evidence of spot modulation (von Essen et al., 2019). Despite our data revealing the strongest temporal variations among G430L visits across the entire sample, the star appears to be quiet according to the rest of the stellar activity metrics we defined.
@@ F.7 HD 209458 b
Stellar Activity: Reported S-indexes () and values: Knutson et al. (2010) report an S-index of = 0.160 and corresponding = -4.970. Sreejith et al. (2020) report a value of = -4.920.029. Claudi et al. (2024) report an S-index of = 0.1560.003 and a corresponding = -4.96850.0210, or = -4.85780.0261 when the correction detailed in Mittag et al. (2013) is applied. Literature values are consistent with each other and indicative of a low level of stellar activity for HD 209458. All of our activity indicators prefer a quiet star over an active one.
@@ F.8 WASP-17 b
Chemistry Results: We find reasonable evidence for clouds in both scenarios but the retrieved values are not consistent between visits (() = 3.79 Pa (Case 1) and () = 1.35 Pa (Case 2)). No detection of K in either Case.
@@ F.10 WASP-31 b
Stellar Activity: Reported S-indexes () and values: Sing et al. (2016) state a value of = -5.225. Danielski et al. (2022) find an S-index = 0.150.00 but they note that WASP-31 is outside of the magnitude range required for their self-consistent conversion to and therefore do not calculate this. Claudi et al. (2024) find an S-index of = 0.1500.023 and a corresponding = -4.98730.1462, or = -4.92220.2079 when the correction detailed in Mittag et al. (2013) is applied. All of the reported values in the literature are consistent with a low level of activity for WASP-31. Bayes Factor and SAD indicate absent-to-low levels of contamination in addition to a small SAT value.
@@ F.12 WASP-69 b
Stellar Activity: Reported S-indexes () and values: Sreejith et al. (2020) report a = -4.540.038. Claudi et al. (2024) report an S-index of = 0.6710.013 and a corresponding = -4.52210.0082, or = -4.16250.0089 when the correction detailed in Mittag et al. (2013) is applied. These reported values are consistent with a moderate-to-high level of activity for WASP-69. The Bayes Factor prefers the base model over the stellar model in both scenarios, however the inclusion of clouds strongly drives the preference towards the stellar activity model. Overall we find multiple indications for a quiet star.
@@ F.14 WASP-80 b
Stellar Activity: Reported S-indexes () and values: Claudi et al. (2024) report an S-index of = 1.7730.087 and a corresponding = -4.82420.0210, or = -4.17070.00213 when the correction detailed in Mittag et al. (2013) is applied. There is a large difference between the derived values depending on whether or not the correction is used making it harder to categorise the expected activity level of WASP-80. Taking the value derived solely with the Noyes et al. (1984) method indicates a low-to-moderate level of activity, whereas, following the Mittag et al. (2013) correction, the implied activity level is much higher. WASP-80 is a potentially active K7V dwarf (Triaud et al., 2013) with a strong chromospheric activity (log R=-4.495). Despite this, the non-detection of stellar spot crossings implies that, if present, spots must be small and symmetrically distributed on the stellar surface (Mancini et al., 2014). Our findings are in agreement with a low level of activity for WASP-80, confirmed by a negative Bayes Factor, a maximum SAD of 1.2410.516 and a low SAT in the range of 3.4%.
@@ F.17 WASP-127 b
Stellar Activity: Reported S-indexes () and values: Danielski et al. (2022) find an S-index of = 0.140.00 and a corresponding = -5.120.09. Claudi et al. (2024) report an S-index of = 0.1560.004 and a corresponding = -5.01570.0280 or = -4.79140.0271 when the correction detailed in Mittag et al. (2013) is applied. These literature values are both indicative of a lower level of activity for WASP-127. Our Bayes Factor indicates an inconclusive preference for stellar contamination, coupled with the distance metric which is the lowest among the whole sample. However, a stronger preference for the active star model (Bayes Factor = 7.75) is obtained when considering gray clouds and alkali metals in the retrieval setup, suggesting the potential presence of stellar activity.

### v3

@@ F.1 GJ 436 b
Stellar Activity: Reported S-indexes () and values: a value of = 0.620 and a corresponding = -5.298 were reported in Knutson et al. (2010), Sreejith et al. (2020) report a = -5.0900.001 and most recently, Claudi et al. (2024) report an S-index of = 0.70.3 and a corresponding = -5.330.03. Considering the Ca H&K lines in isolation, all of the reported values in the literature are consistent and indicative of a lower level of stellar activity for GJ 436. Multiple observations spanning almost a full proposed stellar activity cycle (7.80.1 yrs), revealed large variations 38%3% in the summed flux of the major FUV emission lines, and a multitude of brief flares (Loyd et al., 2023). Spitzer Space Telescope showed variations in transit depth in the same bandpass across different visits, an indication for occultation of stellar active regions (Knutson et al., 2011). Although our Bayes factor results suggest a low preference for an active star model, the SAD favours a substantial degree of stellar contamination. However, caution in the interpretation of this metric is advised due to the considerable uncertainties associated with the optical data.
@@ F.2 GJ 3470 b
Stellar Activity: Reported S-indexes () and values: Sreejith et al. (2020) report a = -4.86. Claudi et al. (2024) report an S-index of = 1.80.3 and a corresponding = -5.060.08. The inconsistency between these reported measurements could be due to differing activity levels at each epoch, although broadly speaking both of the above literature values are indicative of a low-to-moderate activity regime for GJ 3470. STIS observations as analysed by Benneke et al. (2019) do not display any signs of stellar activity, while our interpretation indicate a low level of activity across the stellar metrics we defined, with substantial spectral temporal variations. The monitoring of flaring activity with the Cosmic Origins Spectrograph (COS) provided the detection of three flares (Bourrier et al., 2021), on top of confirming the absorption signature originating from GJ 3470 b’s exosphere in the stellar Lyman- line.
@@ F.3 HAT-P-1 b
Stellar Activity: Reported S-indexes () and values: In the discovery paper, Bakos et al. (2007a) report a value of = -4.720.05. Knutson et al. (2010) calculate an S-index of = 0.158 and a corresponding = -4.984. Danielski et al. (2022) find an S-index of = 0.15 0.00 and a corresponding = -5.110.07, and Claudi et al. (2024) report an S-index of = 0.150.01 and a corresponding = -5.00.1, or = -4.90.1 when the correction detailed in Mittag et al. (2013) is applied. The literature values are consistent and all imply a low-to-moderate level of stellar activity for HAT-P-1. Our Bayes factor results return two of the highest values across the whole sample, indicating a high level of stellar activity. However, the SAD and SAT indicate a low preference for the stellar contamination model and show only minor temporal deviations among the STIS G430L observations which were taken four months apart.
@@ F.4 HAT-P-12 b
Stellar Activity: Reported S-indexes () and values: Knutson et al. (2010) report an S-index of = 0.253 and a corresponding = -5.104. Observations from two different epochs were analysed by Mancini et al. (2018) yielding values of = 0.360.05 and corresponding = -4.880.07 (14.03.2015) and = 0.380.08 and corresponding = -4.870.09 (24.04.2015) respectively. Claudi et al. (2024) find an S-index of = 0.360.06 and a corresponding = -4.910.08, or = -4.540.09 when the correction detailed in Mittag et al. (2013) is applied. Some variation is observed in the literature values but broadly they are suggestive of low-to-moderate activity for HAT-P-12. Jiang et al. (2021) report that the discrepant chemical results among different transmission studies can be attributed to varying levels of unocculted stellar spots and faculae. Using the GTC OSIRIS observatory centred in the optical (500-900 nm), they infer a photosphere covered by 33% of spots and 9% of faculae. We find an overall quiet star across all our metrics, even though the SAT reported is moderately high within our the sample at 6% 2%, indicating a relatively large temporal variability across 20 days during which the system was observed with STIS G430L.
@@ F.5 HAT-P-18 b
Stellar Activity: Reported S-indexes () and values: Claudi et al. (2024) report an S-index of = 0.330.08 and corresponding = -4.90.1, or = -4.50.2 when the correction detailed in Mittag et al. (2013) is applied. These values are consistent with a moderate-to-high level of activity for HAT-P-12. Recent JWST observations with NIRISS are also suggestive of the presence of stellar contamination (Fournier-Tondreau et al., 2024). Both the Bayes factor and the SAD metric agree to a moderate-to-high level of stellar contamination.
@@ F.6 HAT-P-26 b
Stellar Activity: Reported S-indexes () and values: In the discovery paper, Hartman et al. (2011a) report an S-index of = 0.182 and a corresponding = -4.992. Claudi et al. (2024) determine an S-index of = 0.170.08 and a corresponding = -5.00.2, or = -4.90.3 when the correction detailed in Mittag et al. (2013) is applied. Literature values are consistent with a lower level of stellar activity for HAT-P-26. The photometric follow-up of the host star did not reveal evidence of spot modulation (von Essen et al., 2019). Despite our data revealing the strongest temporal variations among G430L visits across the entire sample, the star appears to be quiet according to the rest of the stellar activity metrics we defined.
@@ F.7 HD 209458 b
Stellar Activity: Reported S-indexes () and values: Knutson et al. (2010) report an S-index of = 0.160 and corresponding = -4.970. Sreejith et al. (2020) report a value of = -4.920.03. Claudi et al. (2024) report an S-index of = 0.1560.003 and a corresponding = -4.970.02, or = -4.860.03 when the correction detailed in Mittag et al. (2013) is applied. Literature values are consistent with each other and indicative of a low level of stellar activity for HD 209458. All of our activity indicators prefer a quiet star over an active one.
@@ F.8 WASP-17 b
Chemistry Results: We find reasonable evidence for clouds in both scenarios but the retrieved values are not consistent between visits (() = 3.8 Pa (Case 1) and () = 1.4 Pa (Case 2)). No detection of K in either Case.
@@ F.10 WASP-31 b
Stellar Activity: Reported S-indexes () and values: Sing et al. (2016) state a value of = -5.225. Danielski et al. (2022) find an S-index = 0.150.00 but they note that WASP-31 is outside of the magnitude range required for their self-consistent conversion to and therefore do not calculate this. Claudi et al. (2024) find an S-index of = 0.150.02 and a corresponding = -5.00.1, or = -4.90.2 when the correction detailed in Mittag et al. (2013) is applied. All of the reported values in the literature are consistent with a low level of activity for WASP-31. Bayes factor and SAD indicate absent-to-low levels of contamination in addition to a small SAT value.
@@ F.12 WASP-69 b
Stellar Activity: Reported S-indexes () and values: Sreejith et al. (2020) report a = -4.540.04. Claudi et al. (2024) report an S-index of = 0.670.01 and a corresponding = -4.5220.008, or = -4.1630.009 when the correction detailed in Mittag et al. (2013) is applied. These reported values are consistent with a moderate-to-high level of activity for WASP-69. The Bayes factor prefers the base model over the stellar model in both scenarios, however the inclusion of clouds strongly drives the preference towards the stellar activity model. Overall we find multiple indications for a quiet star.
@@ F.14 WASP-80 b
Stellar Activity: Reported S-indexes () and values: Claudi et al. (2024) report an S-index of = 1.770.09 and a corresponding = -4.820.02, or = -4.1710.002 when the correction detailed in Mittag et al. (2013) is applied. There is a large difference between the derived values depending on whether or not the correction is used making it harder to categorise the expected activity level of WASP-80. Taking the value derived solely with the Noyes et al. (1984) method indicates a low-to-moderate level of activity, whereas, following the Mittag et al. (2013) correction, the implied activity level is much higher. WASP-80 is a potentially active K7V dwarf (Triaud et al., 2013) with a strong chromospheric activity (log R=-4.495). Despite this, the non-detection of stellar spot crossings implies that, if present, spots must be small and symmetrically distributed on the stellar surface (Mancini et al., 2014). Our findings are in agreement with a low level of activity for WASP-80, confirmed by a negative Bayes factor, a maximum SAD of 1.20.5 and a low SAT in the range of 3.4%.
@@ F.17 WASP-127 b
Stellar Activity: Reported S-indexes () and values: Danielski et al. (2022) find an S-index of = 0.140.00 and a corresponding = -5.120.09. Claudi et al. (2024) report an S-index of = 0.1560.004 and a corresponding = -5.020.03 or = -4.790.03 when the correction detailed in Mittag et al. (2013) is applied. These literature values are both indicative of a lower level of activity for WASP-127. Our Bayes factor indicates an inconclusive preference for stellar contamination, coupled with the distance metric which is the lowest among the whole sample. However, a stronger preference for the active star model (Bayes factor = 7.75) is obtained when considering gray clouds and alkali metals in the retrieval setup, suggesting the potential presence of stellar activity.